# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

### v1

@@ IV-C1 Effect of Key Modules
Airport | 27.43 | 0.7387 | 28.48 | 0.7765 | 28.63 | 0.7823 | 28.69 | 0.7837 | 28.72 | 0.7856 | 28.59 | 0.7805 | 28.68 | 0.7851 | 28.72 | 0.7853 |
Bare Land | 34.45 | 0.8344 | 34.84 | 0.8422 | 34.92 | 0.8439 | 34.94 | 0.8445 | 34.96 | 0.8449 | 34.91 | 0.8436 | 34.93 | 0.8444 | 34.97 | 0.8449 |
Baseball Field | 29.19 | 0.7786 | 30.27 | 0.8094 | 30.42 | 0.8137 | 30.45 | 0.8144 | 30.52 | 0.8160 | 30.35 | 0.8123 | 30.43 | 0.8147 | 30.53 | 0.8160 |
Beach | 31.32 | 0.7947 | 31.86 | 0.8082 | 31.94 | 0.8103 | 31.99 | 0.8117 | 32.00 | 0.8118 | 31.94 | 0.8096 | 31.95 | 0.8113 | 32.00 | 0.8115 |
Bridge | 30.19 | 0.8063 | 31.27 | 0.8312 | 31.46 | 0.8353 | 31.52 | 0.8363 | 31.56 | 0.8372 | 31.42 | 0.8339 | 31.52 | 0.8368 | 31.57 | 0.8371 |
Center | 25.76 | 0.6980 | 27.18 | 0.7590 | 27.38 | 0.7665 | 27.47 | 0.7694 | 27.51 | 0.7706 | 27.33 | 0.7645 | 27.43 | 0.7692 | 27.51 | 0.7705 |
Church | 23.09 | 0.6084 | 24.41 | 0.6807 | 24.56 | 0.6895 | 24.61 | 0.6920 | 24.65 | 0.6941 | 24.52 | 0.6875 | 24.56 | 0.6919 | 24.67 | 0.6947 |
Commercial | 26.72 | 0.7178 | 27.59 | 0.7582 | 27.73 | 0.7645 | 27.77 | 0.7660 | 27.81 | 0.7679 | 27.68 | 0.7623 | 27.72 | 0.7659 | 27.81 | 0.7681 |
D-Residential | 23.45 | 0.6301 | 24.46 | 0.6922 | 24.60 | 0.7004 | 24.64 | 0.7026 | 24.69 | 0.7051 | 24.56 | 0.6985 | 24.60 | 0.7029 | 24.70 | 0.7056 |
Desert | 37.14 | 0.8869 | 37.54 | 0.8933 | 37.58 | 0.8944 | 37.63 | 0.8951 | 37.62 | 0.8949 | 37.58 | 0.8941 | 37.60 | 0.8946 | 37.62 | 0.8950 |
Farmland | 32.32 | 0.7934 | 33.18 | 0.8166 | 33.33 | 0.8212 | 33.35 | 0.8216 | 33.42 | 0.8236 | 33.30 | 0.8202 | 33.39 | 0.8235 | 33.43 | 0.8240 |
Forest | 27.67 | 0.6411 | 27.84 | 0.6514 | 27.90 | 0.6558 | 27.95 | 0.6605 | 27.95 | 0.6604 | 27.88 | 0.6552 | 27.92 | 0.6598 | 27.97 | 0.6616 |
Industrial | 26.29 | 0.7063 | 27.43 | 0.7566 | 27.62 | 0.7647 | 27.67 | 0.7669 | 27.71 | 0.7689 | 27.56 | 0.7621 | 27.62 | 0.7665 | 27.72 | 0.7690 |
Meadow | 32.07 | 0.7100 | 32.33 | 0.7136 | 32.38 | 0.7157 | 32.40 | 0.7167 | 32.41 | 0.7172 | 32.37 | 0.7154 | 32.39 | 0.7175 | 32.42 | 0.7172 |
M-Residential | 26.32 | 0.6578 | 27.40 | 0.7055 | 27.55 | 0.7124 | 27.58 | 0.7136 | 27.63 | 0.7159 | 27.52 | 0.7113 | 27.55 | 0.7140 | 27.64 | 0.7164 |
Mountain | 28.05 | 0.6978 | 28.41 | 0.7127 | 28.48 | 0.7163 | 28.50 | 0.7184 | 28.50 | 0.7183 | 28.46 | 0.7152 | 28.45 | 0.7171 | 28.50 | 0.7183 |
Park | 27.14 | 0.7060 | 27.77 | 0.7332 | 27.87 | 0.7383 | 27.91 | 0.7406 | 27.93 | 0.7415 | 27.83 | 0.7365 | 27.86 | 0.7395 | 27.94 | 0.7422 |
Parking | 24.11 | 0.7310 | 26.18 | 0.8059 | 26.51 | 0.8159 | 26.68 | 0.8205 | 26.76 | 0.8225 | 26.46 | 0.8145 | 26.70 | 0.8226 | 26.77 | 0.8228 |
Playground | 29.33 | 0.7723 | 30.70 | 0.8132 | 30.94 | 0.8200 | 31.03 | 0.8216 | 31.10 | 0.8235 | 30.88 | 0.8182 | 31.03 | 0.8230 | 31.12 | 0.8237 |
Pond | 28.76 | 0.7559 | 29.38 | 0.7749 | 29.46 | 0.7781 | 29.49 | 0.7790 | 29.51 | 0.7796 | 29.44 | 0.7769 | 29.46 | 0.7788 | 29.51 | 0.7796 |
Port | 25.83 | 0.7727 | 26.94 | 0.8207 | 27.11 | 0.8255 | 27.18 | 0.8273 | 27.21 | 0.8283 | 27.05 | 0.8237 | 27.13 | 0.8267 | 27.21 | 0.8278 |
Railway Station | 27.00 | 0.6965 | 27.87 | 0.7347 | 28.04 | 0.7434 | 28.11 | 0.7456 | 28.13 | 0.7474 | 27.99 | 0.7404 | 28.07 | 0.7461 | 28.14 | 0.7477 |
Resort | 26.78 | 0.7160 | 27.67 | 0.7547 | 27.81 | 0.7601 | 27.86 | 0.7622 | 27.89 | 0.7634 | 27.78 | 0.7588 | 27.82 | 0.7619 | 27.91 | 0.7638 |
River | 29.29 | 0.7201 | 29.74 | 0.7378 | 29.81 | 0.7412 | 29.84 | 0.7426 | 29.85 | 0.7430 | 29.79 | 0.7401 | 29.80 | 0.7422 | 29.86 | 0.7431 |
School | 25.81 | 0.7046 | 26.81 | 0.7503 | 26.96 | 0.7568 | 27.00 | 0.7587 | 27.04 | 0.7604 | 26.91 | 0.7546 | 26.94 | 0.7577 | 27.05 | 0.7607 |
S-Residential | 25.71 | 0.5790 | 26.15 | 0.6001 | 26.22 | 0.6055 | 26.25 | 0.6070 | 26.27 | 0.6079 | 26.21 | 0.6046 | 26.23 | 0.6075 | 26.27 | 0.6077 |
Square | 27.51 | 0.7286 | 28.71 | 0.7782 | 28.90 | 0.7847 | 28.96 | 0.7870 | 28.99 | 0.7881 | 28.85 | 0.7831 | 28.92 | 0.7868 | 29.01 | 0.7882 |
Stadium | 26.23 | 0.7241 | 27.55 | 0.7771 | 27.75 | 0.7845 | 27.82 | 0.7873 | 27.85 | 0.7882 | 27.68 | 0.7823 | 27.78 | 0.7867 | 27.85 | 0.7881 |
Storage Tanks | 25.15 | 0.6667 | 26.25 | 0.7190 | 26.39 | 0.7255 | 26.41 | 0.7264 | 26.45 | 0.7287 | 26.33 | 0.7233 | 26.39 | 0.7274 | 26.45 | 0.7286 |
Viaduct | 26.77 | 0.6817 | 27.70 | 0.7225 | 27.88 | 0.7310 | 27.91 | 0.7325 | 27.98 | 0.7361 | 27.81 | 0.7277 | 27.87 | 0.7326 | 27.98 | 0.7363 |
Average | 27.89 | 0.7238 | 28.80 | 0.7597 | 28.94 | 0.7653 | 28.99 | 0.7671 | 29.02 | 0.7684 | 28.90 | 0.7637 | 28.96 | 0.7672 | 29.03 | 0.7686 |
IV-C2 Effect of Wavelet Loss
To evaluate the effectiveness of incorporating Wavelet Loss in training, ablation experiments were conducted comparing models trained with Wavelet Loss to those using only Loss. These experiments spanned multiple datasets and degradation models, employing a variety of architectures. The results consistently demonstrated that Wavelet Loss improves PSNR, with SSIM also showing notable enhancement in most cases. On the RSSRD-QH dataset, a decrease in SSIM performance was observed, which was attributed to the complexity of basin images, characterized by intricate textures and diverse structures. Wavelet Loss emphasized high-frequency details, which didn’t align with SSIM’s focus on global structural similarity. A decline in LPIPS was also observed across all experiments with Wavelet Loss, due to its emphasis on high-frequency details and structural accuracy. The prioritization of fine-grained features introduced perceptual differences, as captured by LPIPS. Nonetheless, the substantial improvements in PSNR and SSIM underscored the overall effectiveness of Wavelet Loss in producing high-quality SR images, guiding the correct reconstruction of textures across various architectures.
As shown in Figure 9, despite MambaIR’s global receptive field and its superior performance in experiments, the model still struggled to accurately reconstruct certain textures without Wavelet Loss. The inclusion of Wavelet Loss significantly improved the reconstruction, particularly in the high-frequency components (LH, HL, HH), where fine-grained details were preserved. This highlights Wavelet Loss’s role in guiding the model to recover intricate structures and accurately reconstruct complex textures, thereby enhancing overall image quality.
@@ IV-D1 Comparative Methods
CNN-based models, including SRCNN [6] and EDSR [9]; GAN-based models, including SRGAN [8]; Transformer-based models, including SwinIR [14] and HAT [15]; CNN-Transformer hybrid models, including ConvFormerSR [16]; and Mamba-based models, including MambaIR [18] and FreMamba [20], were involved in evaluating the SR performance of our GDSR against SOTA methods on RSI-SR. We also report the results of our model variants, GDSR_TC and GDSR_MC, as well as the large-scale variant GDSR_L.
@@ IV-D2 Quantitative Evaluations
The quantitative results on the AID and RSSRD-QH datasets are presented in Table V. Our proposed GDSR and GDSR_L models achieved superior performance across all metrics, demonstrating their remarkable SR capabilities in diverse remote sensing tests. On the AID and RSSRD-QH datasets, GDSR achieved performance comparable to MambaIR. On the AID_CDM and RSSRD-QH_CDM datasets, GDSR’s advantages became more pronounced. It achieved PSNR, SSIM, and LPIPS values of 24.10, 0.5770, and 0.5986, respectively, on AID_CDM, and PSNR and SSIM values of 22.86 and 0.3861, respectively, on RSSRD-QH_CDM, significantly outperforming other methods. GDSR_L consistently led across all tests, and both GDSR_TC and GDSR_MC achieved performance comparable to FreMamba on the AID and RSSRD-QH datasets while demonstrating substantial improvements on the AID_CDM and RSSRD-QH_CDM datasets. Specifically, GDSR_MC achieved a 0.14 dB PSNR improvement on AID_CDM and a 0.24 dB improvement on RSSRD-QH_CDM compared to FreMamba.
@@
To further evaluate the generalization capability of the super-resolution models across diverse remote sensing scenarios, we report the PSNR and SSIM results across 30 scene categories in the AID dataset, as shown in Table VII. As evidenced, GDSR exhibited stronger generalization capability compared to state-of-the-art methods, achieving the best performance in nearly all remote sensing scenarios. Specifically, for the ”Parking” and ”Playground” scenes, GDSR outperformed HAT by 0.09 dB in PSNR and by 0.0023 and 0.0021 in SSIM, respectively. Additionally, in the most complex categories, such as ”Church” and ”DenseResidential,” GDSR also achieved the best performance.
@@
Moreover, as the receptive field of the models increased, performance trended upward for CNN-based EDSR, Transformer-based SwinIR, and Mamba-based MambaIR. This aligned with our experimental results on the RGEG Branch and RDEG Branch, where larger receptive fields enabled the activation of more pixels for image reconstruction, enhancing performance. However, the results from MambaIR and FreMamba, along with the experiments on F2B and F2C, indicated that spatial global modeling in complex remote sensing images had reached a bottleneck. Simply increasing the receptive field no longer yielded significant performance gains. A more practical solution was to effectively utilize large receptive fields to activate the appropriate regions and pixels for accurate image reconstruction. The outstanding generalization capability and superior performance of GDSR aligned with our motivation to address the challenges of large-scale remote sensing images efficiently through the introduction of a global-detail dual-branch structure.
@@ IV-D3 Complexity and Efficiency Evaluation
To evaluate the complexity and computational efficiency of our proposed models, we conducted quantitative assessments across various metrics. As shown in Table VIII, the bicubic interpolation technique inherently avoids the need for parameter optimization, resulting in the fastest inference speed within our evaluation scope. The CNN-based model achieved the highest inference speed among all tested deep learning models, primarily due to the inherent efficiency of convolutional parallel operations. The EDSR model, with its deep residual structure, had the largest number of parameters and FLOPs, yet achieved inference speeds several times faster than other architectures.
Transformer-based models, such as SwinIR and HAT, exhibited slower inference speeds compared to CNN-based models but achieved significant performance gains. Models based on Mamba, including MambaIR and FreMamba, leveraged their linear complexity and optimized inference processes, delivering superior performance along with relatively faster inference speeds. ConvFormerSR, which integrated Swin Transformer with CNN architectures, achieved faster inference speeds while maintaining competitive performance.
GDSR, by incorporating RWKV with CNN, not only reduced the parameter count but also achieved inference speeds several times faster than Transformer- and Mamba-based models, while delivering enhanced performance. GDSR_TC and GDSR_MC further optimized performance by combining Swin Transformer and Mamba with CNN, leading to a reduction in parameter size and improved inference speed, without sacrificing performance. GDSR_L, which increased the network depth, consistently outperformed other models in experimental evaluations while maintaining reasonable inference speeds. As demonstrated in Tables II, III, and IV, these results underscore the superiority of the Global-Detail dual-branch architecture design.
Comparing Tables VIII and V reveals that the complexity of our proposed model is comparable to most state-of-the-art models, while its efficiency is superior, and its performance is significantly better. These results highlight the practical potential of our model, demonstrating an effective balance between computational efficiency and performance, making it well-suited for real-world applications.
Model | #Param. | FLOPs | FPS |
Bicubic | - | - | 21505.6 |
SRCNN [6] | 0.02M | 10.47G | 533.1 |
SRGAN [8] | 1.55M | 107.68G | 151.9 |
EDSR [9] | 43.68M | 2519.95G | 15.5 |
SwinIR [14] | 16.62M | 1005.01G | 2.3 |
HAT [15] | 20.81M | 1467.38G | 2.2 |
MambaIR [18] | 15.11M | 838.91G | 2.8 |
ConvFormerSR [16] | 16.49M | 994.90G | 3.2 |
FreMamba [20] | 11.46M | 516.21G | 2.5 |
GDSR_TC (Ours) | 12.83M | 758.82G | 3.3 |
GDSR_MC (Ours) | 13.99M | 788.28G | 4.1 |
GDSR (Ours) | 13.17M | 762.14G | 6.4 |
GDSR_L (Ours) | 25.86M | 1496.14G | 3.2 |

### v4

@@ IV-C2 Effect of Wavelet Loss
| Airport | 27.43 | 0.7387 | 28.48 | 0.7765 | 28.63 | 0.7823 | 28.69 | 0.7837 | 28.72 | 0.7856 | 28.59 | 0.7805 | 28.68 | 0.7851 | 28.74 | 0.7865 |
| Bare Land | 34.45 | 0.8344 | 34.84 | 0.8422 | 34.92 | 0.8439 | 34.94 | 0.8445 | 34.96 | 0.8449 | 34.91 | 0.8436 | 34.93 | 0.8444 | 34.98 | 0.8455 |
| Baseball Field | 29.19 | 0.7786 | 30.27 | 0.8094 | 30.42 | 0.8137 | 30.45 | 0.8144 | 30.52 | 0.8160 | 30.35 | 0.8123 | 30.43 | 0.8147 | 30.55 | 0.8169 |
| Beach | 31.32 | 0.7947 | 31.86 | 0.8082 | 31.94 | 0.8103 | 31.99 | 0.8117 | 32.00 | 0.8118 | 31.94 | 0.8096 | 31.95 | 0.8113 | 32.03 | 0.8124 |
| Bridge | 30.19 | 0.8063 | 31.27 | 0.8312 | 31.46 | 0.8353 | 31.52 | 0.8363 | 31.56 | 0.8372 | 31.42 | 0.8339 | 31.52 | 0.8368 | 31.59 | 0.8378 |
| Center | 25.76 | 0.6980 | 27.18 | 0.7590 | 27.38 | 0.7665 | 27.47 | 0.7694 | 27.51 | 0.7706 | 27.33 | 0.7645 | 27.43 | 0.7692 | 27.52 | 0.7714 |
| Church | 23.09 | 0.6084 | 24.41 | 0.6807 | 24.56 | 0.6895 | 24.61 | 0.6920 | 24.65 | 0.6941 | 24.52 | 0.6875 | 24.56 | 0.6919 | 24.68 | 0.6954 |
| Commercial | 26.72 | 0.7178 | 27.59 | 0.7582 | 27.73 | 0.7645 | 27.77 | 0.7660 | 27.81 | 0.7679 | 27.68 | 0.7623 | 27.72 | 0.7659 | 27.83 | 0.7692 |
| D-Residential | 23.45 | 0.6301 | 24.46 | 0.6922 | 24.60 | 0.7004 | 24.64 | 0.7026 | 24.69 | 0.7051 | 24.56 | 0.6985 | 24.60 | 0.7029 | 24.72 | 0.7067 |
| Desert | 37.14 | 0.8869 | 37.54 | 0.8933 | 37.58 | 0.8944 | 37.63 | 0.8951 | 37.62 | 0.8949 | 37.58 | 0.8941 | 37.60 | 0.8946 | 37.65 | 0.8954 |
| Farmland | 32.32 | 0.7934 | 33.18 | 0.8166 | 33.33 | 0.8212 | 33.35 | 0.8216 | 33.42 | 0.8236 | 33.30 | 0.8202 | 33.39 | 0.8235 | 33.45 | 0.8245 |
| Forest | 27.67 | 0.6411 | 27.84 | 0.6514 | 27.90 | 0.6558 | 27.95 | 0.6605 | 27.95 | 0.6604 | 27.88 | 0.6552 | 27.92 | 0.6598 | 27.99 | 0.6617 |
| Industrial | 26.29 | 0.7063 | 27.43 | 0.7566 | 27.62 | 0.7647 | 27.67 | 0.7669 | 27.71 | 0.7689 | 27.56 | 0.7621 | 27.62 | 0.7665 | 27.74 | 0.7701 |
| Meadow | 32.07 | 0.7100 | 32.33 | 0.7136 | 32.38 | 0.7157 | 32.40 | 0.7167 | 32.41 | 0.7172 | 32.37 | 0.7154 | 32.39 | 0.7175 | 32.43 | 0.7176 |
| M-Residential | 26.32 | 0.6578 | 27.40 | 0.7055 | 27.55 | 0.7124 | 27.58 | 0.7136 | 27.63 | 0.7159 | 27.52 | 0.7113 | 27.55 | 0.7140 | 27.66 | 0.7170 |
| Mountain | 28.05 | 0.6978 | 28.41 | 0.7127 | 28.48 | 0.7163 | 28.50 | 0.7184 | 28.50 | 0.7183 | 28.46 | 0.7152 | 28.45 | 0.7171 | 28.52 | 0.7192 |
| Park | 27.14 | 0.7060 | 27.77 | 0.7332 | 27.87 | 0.7383 | 27.91 | 0.7406 | 27.93 | 0.7415 | 27.83 | 0.7365 | 27.86 | 0.7395 | 27.96 | 0.7431 |
| Parking | 24.11 | 0.7310 | 26.18 | 0.8059 | 26.51 | 0.8159 | 26.68 | 0.8205 | 26.76 | 0.8225 | 26.46 | 0.8145 | 26.70 | 0.8226 | 26.81 | 0.8238 |
| Playground | 29.33 | 0.7723 | 30.70 | 0.8132 | 30.94 | 0.8200 | 31.03 | 0.8216 | 31.10 | 0.8235 | 30.88 | 0.8182 | 31.03 | 0.8230 | 31.13 | 0.8244 |
| Pond | 28.76 | 0.7559 | 29.38 | 0.7749 | 29.46 | 0.7781 | 29.49 | 0.7790 | 29.51 | 0.7796 | 29.44 | 0.7769 | 29.46 | 0.7788 | 29.53 | 0.7802 |
| Port | 25.83 | 0.7727 | 26.94 | 0.8207 | 27.11 | 0.8255 | 27.18 | 0.8273 | 27.21 | 0.8283 | 27.05 | 0.8237 | 27.13 | 0.8267 | 27.22 | 0.8280 |
| Railway Station | 27.00 | 0.6965 | 27.87 | 0.7347 | 28.04 | 0.7434 | 28.11 | 0.7456 | 28.13 | 0.7474 | 27.99 | 0.7404 | 28.07 | 0.7461 | 28.16 | 0.7489 |
| Resort | 26.78 | 0.7160 | 27.67 | 0.7547 | 27.81 | 0.7601 | 27.86 | 0.7622 | 27.89 | 0.7634 | 27.78 | 0.7588 | 27.82 | 0.7619 | 27.92 | 0.7644 |
| River | 29.29 | 0.7201 | 29.74 | 0.7378 | 29.81 | 0.7412 | 29.84 | 0.7426 | 29.85 | 0.7430 | 29.79 | 0.7401 | 29.80 | 0.7422 | 29.88 | 0.7439 |
| School | 25.81 | 0.7046 | 26.81 | 0.7503 | 26.96 | 0.7568 | 27.00 | 0.7587 | 27.04 | 0.7604 | 26.91 | 0.7546 | 26.94 | 0.7577 | 27.07 | 0.7619 |
| S-Residential | 25.71 | 0.5790 | 26.15 | 0.6001 | 26.22 | 0.6055 | 26.25 | 0.6070 | 26.27 | 0.6079 | 26.21 | 0.6046 | 26.23 | 0.6075 | 26.29 | 0.6080 |
| Square | 27.51 | 0.7286 | 28.71 | 0.7782 | 28.90 | 0.7847 | 28.96 | 0.7870 | 28.99 | 0.7881 | 28.85 | 0.7831 | 28.92 | 0.7868 | 29.03 | 0.7891 |
| Stadium | 26.23 | 0.7241 | 27.55 | 0.7771 | 27.75 | 0.7845 | 27.82 | 0.7873 | 27.85 | 0.7882 | 27.68 | 0.7823 | 27.78 | 0.7867 | 27.87 | 0.7889 |
| Storage Tanks | 25.15 | 0.6667 | 26.25 | 0.7190 | 26.39 | 0.7255 | 26.41 | 0.7264 | 26.45 | 0.7287 | 26.33 | 0.7233 | 26.39 | 0.7274 | 26.47 | 0.7298 |
| Viaduct | 26.77 | 0.6817 | 27.70 | 0.7225 | 27.88 | 0.7310 | 27.91 | 0.7325 | 27.98 | 0.7361 | 27.81 | 0.7277 | 27.87 | 0.7326 | 28.00 | 0.7378 |
| Average | 27.89 | 0.7238 | 28.80 | 0.7597 | 28.94 | 0.7653 | 28.99 | 0.7671 | 29.02 | 0.7684 | 28.90 | 0.7637 | 28.96 | 0.7672 | 29.05 | 0.7694 |
@@ IV-D1 Comparative Methods
CNN-based models, including SRCNN [6], EDSR [9], and HAUNet [25]; GAN-based models, including SRGAN [8]; Transformer-based models, including SwinIR [12], HAT [13], and SPT [27]; CNN-Transformer hybrid models, including ConvFormerSR [28]; and Mamba-based models, including MambaIR [15] and FreMamba [16], were involved in evaluating the SR performance of our GDSR against SOTA methods on RSI-SR.
@@ IV-D2 Quantitative Evaluations
The quantitative evaluation results across different datasets and scaling factors were comprehensively analyzed through Tables VII and VIII to assess reconstruction accuracy and perceptual quality. On the AID dataset under bicubic degradation, the proposed GDSR demonstrated state-of-the-art performance with PSNR values of 31.81 dB (×2), 29.05 dB (×3), and 27.39 dB (×4), showing improvements of 0.04-0.11 dB over transformer-based HAT and CNN-based EDSR. The structural preservation capability was further verified by SSIM metrics, where GDSR achieved 0.8592 (×2), 0.7694 (×3), and 0.7022 (×4). Although FreMamba exhibited superior LPIPS scores across ×2-4 scales, GDSR maintained competitive perceptual quality while balancing pixel-wise accuracy and visual realism. For more challenging CDM-synthesized data, GDSR showed remarkable robustness with consistent performance gains, surpassing MambaIR by 0.11 dB PSNR and 0.26% SSIM in ×2 super-resolution. The superiority became more evident on the UCMerced dataset, where GDSR established new benchmarks in 7 out of 12 metrics. Notably, it achieved 0.17 dB PSNR improvement over MambaIR and 0.11 dB gain against HAT in ×3 super-resolution under CDM degradation. Evaluation on the real-world RSSRD-QH dataset further validated the practical effectiveness of GDSR. Under bicubic degradation, the method attained peak PSNR values of 29.87 dB (×2), 26.73 dB (×3), and 25.12 dB (×4). In CDM degradation scenarios, GDSR exhibited particularly strong performance at high super-resolution factors, reaching 22.87 dB PSNR (×3) and 22.27 dB PSNR (×4) with corresponding SSIM values of 0.3864 and 0.3488.
@@
To further evaluate the generalization capability of the super-resolution models across diverse remote sensing scenarios, we report the PSNR and SSIM results across 30 scene categories in the AID dataset under bicubic degradation, as shown in Table IX. As evidenced, GDSR exhibited stronger generalization capability compared to state-of-the-art methods, achieving the best performance in nearly all remote sensing scenarios. Specifically, for the ”Parking” and ”Playground” scenes, GDSR outperformed HAT by 0.13 dB and 0.1 dB in PSNR and by 0.33% and 0.28% in SSIM, respectively. Additionally, in the most complex categories, such as ”Church” and ”DenseResidential,” GDSR also achieved the best performance.
@@
Moreover, as the receptive field of the models increased, performance trended upward for CNN-based EDSR, Transformer-based SwinIR, and Mamba-based MambaIR. This aligned with our experimental results on the RGEG Branch and RDEG Branch, where larger receptive fields enabled the activation of more pixels for image reconstruction, enhancing performance. However, the results from MambaIR and FreMamba, along with the experiments on F2B and F2C, indicated that spatial global modeling in complex RSIs had reached a bottleneck. Simply increasing the receptive field no longer yielded significant performance gains. A more practical solution was to effectively utilize large receptive fields to activate the appropriate regions and pixels for accurate image reconstruction. The outstanding generalization capability and superior performance of GDSR align with our motivation to address the challenges of large-scale RSIs efficiently through the introduction of a global-detail dual-branch structure.
@@ IV-D3 Complexity and Efficiency Evaluation
To evaluate the complexity and computational efficiency of our proposed models, we conducted quantitative assessments across various metrics. As shown in Table X, the bicubic interpolation technique inherently avoids the need for parameter optimization, resulting in the fastest inference speed within our evaluation scope. The CNN-based model achieved the highest inference speed among all tested deep learning models, primarily due to the inherent efficiency of convolutional parallel operations. The EDSR model, with its deep residual structure, had the largest number of parameters and FLOPs, yet achieved inference speeds several times faster than other architectures.
| Model | #Param. | FLOPs | Memory | FPS |
| Bicubic | - | - | 22MB | 27498.0 |
| SRCNN [6] | 0.02M | 4.65G | 140MB | 1109.0 |
| SRGAN [8] | 1.55M | 47.86G | 300MB | 323.3 |
| EDSR [9] | 43.68M | 1119.98G | 1068MB | 35.1 |
| SwinIR [12] | 16.62M | 446.67G | 516MB | 5.5 |
| HAT [13] | 20.81M | 652.17G | 1446MB | 4.6 |
| HAUNet [25] | 2.64M | 1353.07G | 7658MB | 3.1 |
| SPT [27] | 3.25M | 668.46G | 9974MB | 2.2 |
| ConvFormerSR [28] | 16.49M | 442.18G | 500MB | 7.0 |
| MambaIR [15] | 15.11M | 372.85G | 556MB | 6.4 |
| FreMamba [16] | 11.46M | 229.50G | 546MB | 5.5 |
| GDSR (Ours) | 13.17M | 338.73G | 518MB | 14.8 |
Transformer-based models, such as SwinIR and HAT, exhibited slower inference speeds compared to CNN-based models but achieved significant performance gains. Models based on Mamba, including MambaIR and FreMamba, leveraged their linear complexity and optimized inference processes, delivering superior performance along with relatively faster inference speeds. ConvFormerSR, which integrated Swin Transformer with CNN architectures, achieved faster inference speeds while maintaining competitive performance. GDSR, by incorporating RWKV with CNN, not only reduced the parameter count but also achieved inference speeds several times faster than Transformer- and Mamba-based models, while delivering enhanced performance. As demonstrated in Tables II, III, and V, these results underscore the superiority of the Global-Detail dual-branch architecture design.
Comparing Tables X and VII reveals that the complexity of our proposed model is comparable to most state-of-the-art models, while its efficiency is superior, and its performance is significantly better. These results highlight the practical potential of our model, demonstrating an effective balance between computational efficiency and performance, making it well-suited for real-world applications.